Question 8:

Formaldehyde, CH2O, is a toxic gas with a pungent odour. It is used in large quantities for the manufacture of plastics but is also used for the preservation of biological specimens. Calculate the mass percentages of each element in formaldehyde. 9 marks
Answer #1

Answer :

To calculate mass percentage of each element in formaldehyde.

Molar mass of C = 12 g/mol

Molar mass of H = 1 g/mol

Molar mass of O = 16 g/mol

Molar mass of formaldehyde (CH2O) = ( 1 * molar mass of C) + ( 2 * molar mass of H) + ( 1 * molar mass of O)

= ( 1 * 12) + ( 2 * 1 ) + ( 1 * 16)

= 12 + 2 + 16 = 30 g/mol

Mass of total C in CH2O = 12 g

Mass of total H in CH2O = atomicity of H * molar mass of H = 2 * 1 = 2 g

Mass of total O in CH2O = 16 g

Mass percentage of C = ( mass of C in CH2O / total mass of CH2O ) * 100

= ( 12 / 30 ) * 100 = 1200 / 30 = 40 %

Mass percentage of H = ( mass of H in CH2O/ total mass of CH2O ) * 100

= ( 2 / 30 ) * 100 = 200 / 30 = 6.67 %

Mass percent of O = ( mass of O in CH2O/ total mass of CH2O) * 100

= ( 16 / 30 ) * 100

= 1600 / 30 = 53.33 %

Answer :

Mass percent of C = 40 %

Mass percent of H = 6.67 %

Mass percent of O = 53.33 %